Dorsey Company manufactures three products from a common input in a joint processing operation. Joint processing costs up to the split-off point total $360,000 per quarter. For financial reporting purposes, the company allocates these costs to the joint products on the basis of their relative sales value at the split-off point. Unit selling prices and total output at the split-off point are as follows:

Product Selling Price Quarterly
Output
A $ 22.00 per pound 13,400 pounds
B $ 16.00 per pound 20,900 pounds
C $ 28.00 per gallon 4,600 gallons

Each product can be processed further after the split-off point. Additional processing requires no special facilities. The additional processing costs (per quarter) and unit selling prices after further processing are given below:

Product Additional
Processing Costs
Selling
Price
A $ 75,970 $ 27.30 per pound
B $ 109,395 $ 22.30 per pound
C $ 48,260 $ 36.30 per gallon

Required:

1. What is the financial advantage (disadvantage) of further processing each of the three products beyond the split-off point?

2. Based on your analysis in requirement 1, which product or products should be sold at the split-off point and which product or products should be processed further?
Answer #1

Answer:-1)-The financial disadvantages of processing Product A beyond the split-off point is= $4950.

The financial advantages of processing Product B beyond the split-off point is=$22275.

The financial disadvantages of processing Product C beyond the split-off point is=$10080.

Explanation-

Doresy Company
Statement Of Financial advantage (disadvantage)
Product Sale value at Split-Off-Point Sale value of processed product Incremental sale Cost to further process Incremental profit/(loss)
$ $ $ $ $
(a) (b) (c=b-a) (d) (e=c-d)
A 13400 pounds*$22 per pound =294800 13400 pounds*$27.30 per pound =365820 71020 75970 -4950
B 20900 pounds*$16 per pound = 334400 20900 pounds*$22.30 per pound = 466070 131670 109395 22275
C 4600 gallons*$28 per gallon = 128800 4600 gallons*$36.30 per gallon = 166980 38180 48260 -10080

2)-Product A & C should be sold at split-off point and product B should be processed further.